# Determinations of angular stiffness in rotational optical tweezers

### v1

@@ Abstract
Rotational optical tweezers are used to probe the mechanical properties of unknown microsystems. Quantifying the angular trap stiffness is essential for interpreting the rotational dynamics of probe particles. While methods for trap stiffness calibration are well established for translational degrees of freedom, angular trapping has been largely overlooked and is often assumed to behave analogously to translational dynamics. However, rotational and translational motions are sensitive to distinct experimental parameters and offer separate insights. This work covers passive analysis techniques for calibrating the angular trap stiffness and examines the influence of several factors unique to rotational optical tweezers. We show that the parameters of an ancillary measurement beam can be tuned to minimise its influence on angular trapping dynamics, while offering unprecedented improvements for nanoparticle analysis. We also explore the combined effects of shape-induced and material birefringence in spheroidal vaterite probes, and present a framework for assessing hydrodynamic and inertial contributions. These results provide a foundation for characterising rotational optical tweezers independent from translational models.
@@ I Introduction
Optical tweezers (OTs) is a powerful experimental tool widely employed to investigate both fundamental and applied phenomena at micro-to-nanoscopic size scales [Ashkin1986, Jones2015, volpe_roadmap_2023]. Their success stems from the ability to trap microscopic particles and perform precise measurements of position, angle, force, and torque. When OTs are used to extract physical properties of local microenvironments, it is necessary to use a well-characterised probe particle and accurately establish its interaction with the optical potential. Analysis of the probe’s dynamics can be interpreted to accurately attribute the results to the desired phenomena. This is particularly useful in microrheology studies of biological systems [Norregaard2017, Weigand2017, watson_rotational_2022, Vos2024] and quantifying microscopic forces and torques [Block1990, Singer2003, Fallman2004, Inman2010, CatalaCastro2022, Chakraborty2023, hong_optical_2024].
@@
In most trapping experiments, the optical potential that confines the particle in space is treated as harmonic for small displacements from the equilibrium position, $\bm{x_{0}}$. In this regime, the force is proportional to displacements, such that $\bm{F}_{\mathrm{opt}}(t)=-\bm{\kappa}(\bm{x}(t)-\bm{x}_{0})$ where $\bm{\kappa}=(\kappa_{x},\kappa_{y},\kappa_{z})$ is the translational trap stiffness about each axis. Determining $\kappa$ is an important starting point for quantitative optical tweezers measurements and forms a substantial portion of optical tweezers literature, appearing in reviews of established techniques and as a vital detail in the experimental methodology attached to research articles [berg-sorensen_power_2004, Gieseler2021, mirzaei-ghormish_nonlinear_2025]. Consequently, calibrations have become routine, often favouring the equipartition analysis for a simple estimate of trap stiffness or power spectrum analysis for a more robust and reliable estimate [berg-sorensen_power_2004, Gieseler2021].
@@
Rotational measurements with optical tweezers are interested in either the rotation or alignment dynamics of the probes [Friese1998, Bishop2004, Arita2016, roy_simultaneous_2014, Ma2018, Bennett2013, watson_rotational_2022, watson_interrogating_2025]. In the context of alignment, the optical torque creates an angular confining potential that can be quantified by its angular trap stiffness, $\chi$. Rotational measurements are less common with comparatively less published work that addresses characterising the interaction between the angular dynamics of the probe with the optical potential for applications. However, determining $\chi$ is of vital importance for wideband rotational microrheometry and for rotational ballistic measurements [Zhang2017, watson_interrogating_2025].
Calibrations in rotational optical tweezers have previously focussed on accurate torque detection, [pedaci_calibration_2012], or assume that the angular dynamics are exactly analogous to the translational dynamics, so similar approaches can be employed. However, there are significant differences between rotational and translational tweezers experiments and descriptions of the probe dynamics are not strictly interchangeable. The two regimes have different sensitivities for experimental parameters and offer separate insights into systems. For example, translational and rotational hydrodynamics effects differ vastly, which is particularly important in the context of biofluids and biological microsystems that are inherently inhomogeneous [Leach2009, Gibson2019, makuch_diffusion_2020, michalski_rotational_2024].
@@
This work aims to provide a framework for angular trap stiffness calibration about the beam axis. The results presented here are based on a typical rotational optical tweezers system with vaterite microspheres as birefringent probe particles. The mechanism underpinning this system is detailed in Section II, which describes the apparatus, the expected dynamics of the vaterite probe, and the detection method. Whilst our results focus on this specific configuration, the methodology can be extended to other rotational probes and experimental parameters.
@@
The results are presented across three distinct sections. Section III introduces several analysis techniques that are applied to experimental data to demonstrate the straightforward calibration procedure for determining angular stiffness. Section IV adopts a computational approach, calculating the angular trap stiffness directly from the simulated torque profile. This method allows system parameters to be varied to explore the resulting dynamics. Finally, Section V examines, through computational methods, several factors that influence angular trap stiffness and are unique to rotational optical trapping. These include the contribution from a measurement beam, hydrodynamic and inertial effects, and probe morphology. Overall, this study provides a systematic methodology for determining the angular trap stiffness and highlights the unique consequences for the rotational regime.
@@ II.1 Rotational optical tweezers (ROTs) system
Figure 1: (A) Simplified schematic of the rotational optical tweezers system. (B) Circular polarisation configuration used to calibrate for angular displacement. The plot shows the conversion from volts to radians, where the black traces are segments taken from the time trace of the rotating vaterite and the red line is the mean. (C) Linear polarisation configuration and the measurement of the probe’s azimuthal trajectory are used for angular trap-stiffness calibration.
Rotational optical tweezers (ROTs) exploits the conservation of linear momentum for optical trapping while also introducing rotation by considering the total angular momentum of the system. Photons can carry spin angular momentum of $\pm\hbar/\mathrm{photon}$, as well as orbital angular momentum from the spatial distribution of their wavefront. As the light interacts with a particle, a transfer of angular momentum can occur, producing an optical torque and a change to the total angular momentum flux of the light. Rotational optical tweezers offer highly flexible control over particle dynamics, with tunable beam shape and polarisation enabling precise manipulation of torque, alignment, and rotational confinement. However, we restrict our discussion to the transfer of spin angular momentum to a birefringent probe. Fig. 1 shows the rotational optical trapping system used to achieve the desired dynamics.
The optical trap is formed using a $1064\text{\,}\mathrm{nm}$ Nd:YAG laser focussed by a high numerical aperture objective lens (Olympus UPLSAPO60XW, $60\times$, 1.2 NA, water). The optical power is controlled using an acousto-optic modulator (AOM, DTD-274HD6 driven at $27\text{\,}\mathrm{MHz}$). The forward transmitted light is collected by a high NA condenser lens (Olympus U-AA, 1.4 NA, oil). $D_{0-3}$ are photovoltaic detectors (PDA36A-EC, Thorlabs) used to measure the initial power of the beams ($D_{0}$ for IR, $D_{3}$ for He-Ne) and the power of right ($D_{\ignorespaces\penalty 10000\ 1}$) and left ($D_{2}$) circular polarisation components. Half waveplates ($\lambda/2$) and quarter waveplates ($\lambda/4$) are used to control the polarisation state of the trap and the polarisation basis of the detection schemes.
@@
where $P$ is the power of the trapping beam, $\omega$ is the angular frequency of the trapping light, and $\Delta\sigma=\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$. Here, $\sigma$ is equivalent to the degree of circular polarisation of the light. $\sigma$ is measured by detecting the proportion of right and left circularly polarised light from
@@
The optical torque acting on the probe—in our case, a vaterite microsphere—depends on its anisotropy and inhomogeneity. Vaterite microspheres are widely used for rotation experiments, due to their spherical geometry, strong birefringence, and routine synthesis protocol [Vogel2009]. Vaterites are positive uniaxial crystals that—given certain conditions—can undergo spherulite growth where the extraordinary axis of individual crystal follows a rotationally symmetric hyperbolic distribution [Vogel2009]. An optic axis, defined parallel to this symmetry axis, is used to describe the orientation of the probe. Since the torque detection detects torque about the beam axis, only the azimuthal trajectory, $\phi$, is measured.
@@
where the factor of $2$ accounts for the probes mirror symmetry $D_{4}$ and $D_{5}$ are the voltage signals. A half waveplate is positioned before the polarisation beam splitter cube, so when at the equilibrium orientation the angular sensitivity is maximised. We have incorporated a balanced photodetector (HBPR-100M-60K-SI, Femto) to measure the equivalent of $D_{4}-D_{5}$ with the common mode noise filtered out. This is necessary when optimising the angular sensitivity, as in the case of rotational ballistic measurements, [watson_interrogating_2025]. For this work, where that level of sensitivity was not required, we restricted the angular measurement to the separate signals from $D_{4}$ and $D_{5}$.
@@
In a circularly polarised trap, there is a continuous transfer of spin-angular momentum that generates an optical torque independent of its azimuthal orientation. In this configuration, the microsphere rapidly reaches terminal angular velocity, determined by the balance between the optical torque and viscous drag. The latter is given by $T_{\mathrm{drag}}=8\pi\eta a^{3}\Omega$, where $\eta$ is the viscosity, $a$ the radius, and $\Omega$ the rotation rate. Measurements of the constant rotation are used to calibrate Eq. 3, as the entire voltage readout is sampled. The plot in Fig. 1B shows the conversion from the voltage signal to radians. The recorded voltage varies sinusoidal to twice the azimuthal angle of the vaterite, $V\propto\sin(2\theta)$ [Bennett2013]. Fig. 1B overlays segments of the detected voltage from a rotating vaterite and converts these to azimuthal angle, phiphi, in radians. The black traces have been segments between maxima and minima and the red line shows the mean result. This plot demonstrates that slight variation in trajectory as the particle rotates.
@@
When trapped in linearly polarised light, the probe aligns with the polarisation axis of the incident beam and orthogonal to the propagation direction, producing an equilibrium orientation $\phi_{0}$. Thermal motion results in fluctuations of angular position around the equilibrium. With the angle calibration determined from the probe rotating, the angular trajectory is obtained, allowing for the determination of the angular stiffness or for analysis of angular dynamics used to probe a system of interest. An example trajectory of a probe in this configuration is shown in Fig. 1C. The optical torque about the beam axis is given by $T_{\mathrm{opt}}(t)=(\chi/2)\sin(2(\phi(t)-\phi_{0}))$ which for small angles can be treated as $T_{\mathrm{opt}}(t)=-\chi(\phi(t)-\phi_{0})$. This provides a second method to determine the optical torque using the measurement beam, given $\chi$ is known, that matches the polarisation-based torque detection from the trapping beam.
@@
There are two other rotational degrees of freedom that should be considered, suggesting an angular trap stiffness vector of $\bm{\chi}=(\chi_{\rho},\chi_{\theta},\chi_{\phi})$. The first is the roll of the vaterite, which is its rotation about its optic axis. Due to the rotational symmetry, there is no optical torque generated about this axis so $\chi_{\rho}=0$. The second is the elevation of the vaterite, which is the out-of-plane rotation from the plane orthogonal to the beam propagation direction. Since the optic axis remains in this plane regardless of incident polarisation, there is an associated angular stiffness, $\chi_{\theta}$. This degree of freedom is difficult to measure precisely, particularly with vaterite, but has been investigated using other rotational probes [lokesh_realization_2021, Roy2023]. As mentioned earlier, we restrict our analysis to the azimuthal stiffness, $\chi_{\phi}$.
@@ II.2 Computational methods
The computational results presented in this paper use the Optical Tweezers Toolbox [OTT2007] to calculate the interaction between the probe and the light, via the T-matrix method for scattering calculations [Nieminen2011]. Briefly, this involves modelling the electric field of this incident and scattered light as a series expansion of vector spherical wave functions (VSWF), with expansion coefficients $\bm{a}$ and $\bm{p}$ respectively. The T-matrix describes the coupling between $\bm{a}$ and $\bm{p}$ in the presence of a scatterer, such that $\bm{p}=\textbf{T}\bm{a}$, allowing calculation of the scattered field for calculations of force and torque [Crichton2000, Nieminen2011]. The T-matrix is independent of the incident field, enabling repeated calculations in a particle-centred coordinate system, making it ideal for investigating the force and torques at arbitrary positions and orientations. The T-matrix that represents the scattering of vaterite microspheres was formulated using the discrete dipole approximation with point matching [Loke2009]. The volume of the probe is discretised into an array of closely spaced dipoles ($\lambda/20$ typical spacing), with the polarisability of each dipole defined, accounting for the material anisotropy and inhomogeneity, following the internal structure described by Loke2009.
III Determination of stiffness from angular trajectories
Determination of trap stiffness from stochastic trajectories requires an accurate model of the underlying dynamics. For this case, that is a model describing the azimuthal trajectory of the probe when trapped by linearly polarised light. For clarity, we take $\phi(t)$ as the angular displacement from equilibrium (i.e. choose $\phi_{0}=0$). We assume the azimuthal trajectory is independent of other degrees of freedom, and restrict the following analysis to 1D rotational dynamics motion without any coupling.
The equation of motion follows the Einstein-Ornstein-Uhlenbeck theory of rotational Brownian motion, written as the Langevin equation [Bennett2013],

### v2

@@ Abstract
Rotational optical tweezers are used to probe the mechanical properties of unknown microsystems. Quantifying the angular trap stiffness is essential for interpreting the rotational dynamics of probe particles. While methods to determine trap stiffness are well established for translational degrees of freedom, angular trapping is often treated analogously even though rotational and translational motions are sensitive to distinct experimental parameters and offer separate insights. This work details passive analysis techniques for determining the angular trap stiffness within the linear restoring torque model and examines the influence of several factors unique to rotational optical tweezers. We show that the parameters of an ancillary measurement beam can be tuned to minimise its influence on angular trapping dynamics, providing necessary improvements for nanoparticle-scale analysis. We also explore the combined effects of shape-induced and material birefringence in spheroidal vaterite probes, and present a framework for assessing hydrodynamic and inertial contributions.
@@ I Introduction
Optical tweezers (OTs) is a powerful experimental tool widely employed to investigate both fundamental and applied phenomena at micro-to-nanoscopic size scales [2, 18, 41]. Their success stems from their ability to trap microscopic particles and perform precise measurements of position, angle, force, and torque. When OTs are used to extract physical properties of local microenvironments, it is necessary to use a well-characterised probe particle and accurately establish its interaction with the optical potential. Analysis of the probe’s dynamics can be interpreted to accurately attribute the results to the desired phenomena. This is particularly useful in microrheology studies of biological systems [29, 46, 43, 42] and in quantifying microscopic forces and torques [7, 36, 12, 17, 8, 9, 16].
@@
In most trapping experiments, the optical potential that confines the particle in space is treated as harmonic for small displacements from the equilibrium position, $\bm{x_{0}}$. In this regime, the force is proportional to displacements, such that $\bm{F}_{\mathrm{opt}}(t)=-\bm{\kappa}(\bm{x}(t)-\bm{x}_{0})$ where $\bm{\kappa}=(\kappa_{x},\kappa_{y},\kappa_{z})$ is the translational trap stiffness along each axis. Determining $\kappa$ is an important starting point for quantitative optical tweezers measurements and forms a substantial portion of optical tweezers literature, appearing in reviews of established techniques and as a vital detail in the experimental methodology attached to research articles [4, 15, 26]. Consequently, calibrations have become routine, often favouring equipartition analysis for a simple estimate of trap stiffness or power spectrum analysis for a more robust and reliable estimate [4, 15].
@@
When using rotational optical tweezers (ROTs), we are interested in either the rotation or alignment dynamics of the probes and the corresponding torques [13, 6, 19, 31, 3, 34, 1, 23, 39, 43, 44]. Rotational measurements are less common with comparatively few publications that address the interaction between the angular dynamics of the probe with the optical potential for applications. In the context of alignment, the optical torque creates an angular confining potential that can be quantified by its angular trap stiffness, $\chi$, where $\bm{T}_{\mathrm{opt}}(t)=-\bm{\chi}(\bm{\phi}(t)-\bm{\phi}_{0})$ [3]. Determining $\chi$ is of vital importance for wideband rotational microrheometry and for rotational ballistic measurements [47, 44].
Studies of calibration methods in ROTs have been largely focussed on accurate torque detection [32], rather than the alignment dynamics, where methods to determine angular trap stiffness are treated analogously to translational stiffness due to the linear restoring torque. However, descriptions of the rotational and translational trap stiffnesses are not strictly equivalent as the two regimes are sensitive to different experimental parameters. For example, translational and rotational hydrodynamics effects differ vastly. This is particularly important for microrheological studies of biofluids and biological microsystems, as these systems are inherently inhomogeneous [20, 14, 24, 25].
@@
This work aims to provide a framework for angular trap stiffness calibration about the beam axis demonstrating the linear model and how experiment parameters can influence the stiffness. The results presented here are based on a typical rotational optical tweezers system with vaterite microspheres [40] as birefringent probe particles. The mechanism underpinning this system is detailed in Section II, which describes the apparatus, the expected dynamics of the vaterite probe, and the detection method. Whilst our results focus on this specific configuration, the methodology can be extended to other rotational probes and experimental parameters.
@@
The results are presented in three distinct sections. Section III applies several analysis methods to typical experimental data for determining the angular trap stiffness. These methods are based on the linear model and the results validate their appropriate use. We highlight the maximum likelihood estimation method (Eq. 22) as unique to our rotational optical tweezes system without a translational analogue. Section IV adopts a computational approach, calculating the angular trap stiffness directly from the simulated torque profile and investigate the effect of varying system parameters. Finally, Section V examines, using computational methods, several factors that influence angular trap stiffness and are unique to rotational optical trapping. These include the contribution from a measurement beam, hydrodynamic, inertial effects, and probe morphology. Overall, this study provides a systematic methodology for determining the angular trap stiffness and highlightsunique consequences for the rotational regime.
@@ II.1 Rotational optical tweezers (ROTs) system
Figure 1: (A) Simplified schematic of the rotational optical tweezers system. (B) Circular polarisation configuration used to calibrate for angular displacement. he plot shows the conversion from volts to radians, where the black traces are segments taken from the time trace of the rotating vaterite and the red line is the mean. (C) Linear polarisation configuration and the measurement of the probe’s azimuthal trajectory are used for angular trap-stiffness calibration.
ROTs exploits the conservation of linear momentum for optical trapping while also introducing rotation by considering the total angular momentum of the system. Photons can carry spin angular momentum of $\pm\hbar/\mathrm{photon}$, as well as orbital angular momentum from the spatial distribution of their wavefront. As the light interacts with a particle, a transfer of angular momentum can occur, producing an optical torque and a change to the total angular momentum flux of the light. Rotational optical tweezers offer highly flexible control over particle dynamics, with tunable beam shape and polarisation enabling precise manipulation of torque, alignment, and rotational confinement. However, we restrict our discussion to the transfer of spin angular momentum to a birefringent probe. Fig. 1 shows the rotational optical trapping system used to achieve the desired dynamics.
The optical trap is formed using a $1064\text{\,}\mathrm{nm}$ Nd:YAG laser focussed by a high numerical aperture objective lens (Olympus UPLSAPO60XW, $60\times$, 1.2 NA, water). The optical power is controlled using an acousto-optic modulator (AOM, DTD-274HD6 driven at $27\text{\,}\mathrm{MHz}$). The forward transmitted light is collected by a high NA condenser lens (Olympus U-AA, 1.4 NA, oil). $D_{0-3}$ are photovoltaic detectors (PDA36A-EC, Thorlabs) used to measure the initial power of the beams ($D_{0}$ for IR, $D_{3}$ for He-Ne) and the power of right ($D_{1}$) and left ($D_{2}$) circular polarisation components. Half waveplates ($\lambda/2$) and quarter waveplates ($\lambda/4$) are used to control the polarisation state of the trap and the polarisation basis of the detection schemes.
@@
where $P$ is the power of the trapping beam, $\omega$ is the angular frequency of the trapping light, and $\Delta\sigma=\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$ is the change in polarisation equal to the difference between the incoming and outgoing spin angular momentum density. Here, $\sigma$ is equivalent to the degree of circular polarisation of the light. $\sigma$ is measured by detecting the proportion of right and left circularly polarised light from
@@
The optical torque acting on the probe—in our case, a vaterite microsphere—depends on its anisotropy and inhomogeneity. Vaterite microspheres are widely used for rotation experiments, due to their spherical geometry, strong birefringence, and routine synthesis protocol [40]. Vaterites are positive uniaxial crystals that—given certain conditions—can undergo spherulite growth where the extraordinary axis of individual crystal follows a rotationally symmetric hyperbolic distribution [40]. An optic axis, defined parallel to this symmetry axis, is used to describe the orientation of the probe. Since the torque detection detects torque about the beam axis, only the azimuthal trajectory, $\phi$, is measured.
@@
where the factor of $2$ accounts for the probes mirror symmetry, and $D_{4}$ and $D_{5}$ are the voltage signals. A half waveplate is positioned before the polarisation beam splitter cube to maximise the angular sensitivity when it is at the equilibrium orientation. We have incorporated a balanced photodetector (HBPR-100M-60K-SI, Femto) to measure the equivalent of $D_{4}-D_{5}$ with the common mode noise filtered out. This is necessary when optimising the angular sensitivity, as in the case of rotational ballistic measurements, [44]. For this work, where that level of sensitivity was not required, we restricted the angular measurement to the separate signals from $D_{4}$ and $D_{5}$.
@@
In a circularly polarised trap, there is a continuous transfer of spin-angular momentum that generates an optical torque independent of its azimuthal orientation. In this configuration, the microsphere rapidly reaches terminal angular velocity, determined by the balance between the optical torque and viscous drag. The latter is given by $T_{\mathrm{drag}}=8\pi\eta a^{3}\Omega$, where $\eta$ is the viscosity of the medium, $a$ the radius of the probe, and $\Omega$ the rotation rate. Measurements of the constant rotation are used to calibrate Eq. 3, as the entire voltage readout is sampled. The plot in Fig. 1B shows the conversion from the voltage signal to radians. The recorded voltage varies sinusoidal to twice the azimuthal angle of the vaterite, $V\propto\sin(2\theta)$ [3]. Fig. 1B overlays segments of the detected voltage from a rotating vaterite and converts these to azimuthal angle, $\phi$, in radians. The black traces have been segments between maxima and minima and the red line shows the mean result. This plot demonstrates the slight variation in trajectory as the particle rotates.
@@
When trapped in linearly polarised light, the probe aligns with the polarisation axis of the incident beam and is orthogonal to the propagation direction, producing an equilibrium orientation $\phi_{0}$. Thermal motion results in fluctuations of angular position around the equilibrium. With the angle calibration determined from the probe rotating, the angular trajectory is obtained, allowing for the determination of the angular stiffness or for analysis of angular dynamics used to probe a system of interest. An example trajectory of a probe in this configuration is shown in Fig. 1C. The optical torque about the beam axis is given by $T_{\mathrm{opt}}(t)=(\chi/2)\sin(2(\phi(t)-\phi_{0}))$ which for small angles can be treated as $T_{\mathrm{opt}}(t)=-\chi(\phi(t)-\phi_{0})$. This provides a second method to determine the optical torque using the measurement beam, given $\chi$ is known, that matches the polarisation-based torque detection from the trapping beam.
@@
There are two other rotational degrees of freedom that should be considered, suggesting an angular trap stiffness vector of $\bm{\chi}=(\chi_{\rho},\chi_{\theta},\chi_{\phi})$. The first is the roll of the vaterite, which is its rotation about its optic axis. Due to the rotational symmetry, there is no optical torque generated about this axis so $\chi_{\rho}=0$. The second is the elevation of the vaterite, which is the out-of-plane rotation from the plane orthogonal to the beam propagation direction. Since the optic axis remains in this plane regardless of incident polarisation, there is an associated angular stiffness, $\chi_{\theta}$. This degree of freedom is difficult to measure precisely, particularly with vaterite, but has been investigated using other rotational probes [22, 35]. As mentioned earlier, we restrict our analysis to the azimuthal stiffness, $\chi_{\phi}$, which we refer to as $\chi$.
@@ II.2 Computational methods
The computational results presented in this paper use the Optical Tweezers Toolbox [28] to calculate the interaction between the probe and the light, via the T-matrix method for scattering calculations [27]. Briefly, this involves modelling the electric field of this incident and scattered light as a series expansion of vector spherical wave functions (VSWF), with expansion coefficients $\bm{a}$ and $\bm{p}$ respectively. The T-matrix describes the coupling between $\bm{a}$ and $\bm{p}$ in the presence of a scatterer, such that $\bm{p}=\textbf{T}\bm{a}$, allowing calculation of the scattered field for calculations of force and torque [10, 27]. The T-matrix is independent of the incident field, enabling repeated calculations in a particle-centred coordinate system, making it ideal for investigating the force and torques at arbitrary positions and orientations. The T-matrix that represents the scattering of vaterite microspheres was formulated using the discrete dipole approximation with point matching [21]. The volume of the probe is discretised into an array of closely spaced dipoles ($\lambda/20$ typical spacing), with the polarisability of each dipole defined, accounting for the material anisotropy and inhomogeneity, following the internal structure described by Loke et al. [21].
III Passive Methods for Angular Stiffness Determination
The trap stiffness can be determined by analysing the stochastic trajectories and modelling optical interaction as a linear restoring torque. For clarity, we take $\phi(t)$ as the azimuthal angular displacement from equilibrium and choose $\phi_{0}=0$ when $T_{\mathrm{opt}}=0$. We assume the azimuthal trajectory is independent of other degrees of freedom, and we restrict the following analysis to 1D rotational dynamics motion without any coupling.
The equation of motion follows the Einstein-Ornstein-Uhlenbeck theory of rotational Brownian motion, written as the Langevin equation [3],